# Good involutions of conjugation subquandles

### v4

@@ 7.1. Setup for the proof
To prove Theorem 1.4, we pass to a suitable subquandle $\operatorname{Conj}X$ of a conjugation quandle $\operatorname{Conj}K$ and apply Theorem 6.7. Namely, consider the semidirect product
@@
where $\mathbb{Z}/2\mathbb{Z}$ acts on $G\times G$ by swapping the coordinates. In 2021, Bergman [core]*(6.5) showed that the map $\varphi:G\hookrightarrow K$ defined by
$\varphi(g):=(g,g^{-1},1)$
is a quandle embedding of $\operatorname{Core}G$ into $\operatorname{Conj}K$. Henceforth, let $X:=\varphi(G)$ denote the image of $\varphi$, and let $H:=\langle X\rangle$ be the subgroup of $K$ generated by $X$. First, we make two observations.
Observation 7.1.
Every element of $X$ has order $2$ in $H$.
@@
Consider the subgroup $S$ of $K$ defined to be
$S:=\{(g,g^{-1},i)\mid g\in G,\ i\in\{0,1\}\}\cong G\rtimes\mathbb{Z}/2\mathbb{Z}.$
Then $H=S$.
Proof.
Certainly, $X\subset S$, so $H\leq S$. Conversely, for all elements $(g,g^{-1},0)\in S\setminus X$, we have
$(g,g^{-1},0)=(g,g^{-1},1)(1,1,1)\in H,$
so $S\leq H$. ∎
7.1.1. Calculation of the center
In light of Lemma 7.2, we henceforth notate $H$ as the semidirect product $G\rtimes\mathbb{Z}/2\mathbb{Z}$, where $\mathbb{Z}/2\mathbb{Z}$ acts on $G$ by inversion. In particular, the map
(7.1) $\varphi(g)=(g,1)$
is a quandle isomorphism from $\operatorname{Core}G$ to the subquandle $\operatorname{Conj}X$ of $\operatorname{Conj}H$. Hence, to prove Theorem 1.4, it suffices to apply Theorem 6.7 to $\operatorname{Conj}X$; cf. Observation 3.8. To that end, we first compute the center $Z(H)$.
Lemma 7.3.
Suppose that $G$ contains an element $z$ such that $zgz^{-1}=g^{-1}$ for all $g\in G$. Then $G$ is an elementary abelian $2$-group.
@@ 7.1.1. Calculation of the center
We show that every element $(z,i)$ of $Z(H)$ has the desired form; the reverse implication is straightforward. If $i=1$, then $(z,1)^{-1}=(z,1)$, so for all $g\in G$, the equation
$(g,0)=(z,1)(g,0)(z,1)^{-1}=(zg^{-1}z^{-1},0)$
holds in $H$. It follows from Lemma 7.3 that $G$ is an elementary abelian $2$-group, which contradicts the assumption we made at the beginning of this section.
Thus, we must have $i=0$, so $(z,0)^{-1}=(z^{-1},0)$. It follows that
$(g,0)=(z,0)(g,0)(z,0)^{-1}=(zgz^{-1},0)$
for all $g\in G$, so $z\in Z(G)$. On the other hand,
$(1,1)=(z,0)(1,1)(z,0)^{-1}=(z,1)(z^{-1},0)=(z^{2},1).$
Hence, $z\in A$. ∎
7.2. Results
The following is Theorem 1.4. We use the notation established in the previous subsection, and we denote the rack structure of $\operatorname{Core}G$ by $s$.
Theorem 7.5.
@@ 7.2. Results
Given any function $\alpha:G\to A$, Lemma 7.2 implies that we can define a function $\overline{\alpha}:X\to H$ by
$\overline{\alpha}(g,1):=(\alpha(g),0).$
By Proposition 7.4, the image of $\overline{\alpha}$ lies in $Z(H)$. Conversely, given any function $\zeta:X\to Z(H)$, define a function $\underline{\zeta}:G\to G$ by sending $g\in G$ to the first coordinate of $\zeta(g,1)\in Z(H)$. By Proposition 7.4, the image of $\underline{\zeta}$ lies in $A$. Hence, it is straightforward to verify that the assignments $\alpha\mapsto\overline{\alpha}$ and $\zeta\mapsto\underline{\zeta}$ are mutually inverse.
Now, recall that the map $\varphi:G\to X$ from (7.1) is a quandle isomorphism from $\operatorname{Core}G$ to $\operatorname{Conj}X$. In light of Observation 3.8, a straightforward calculation of $\varphi\rho\varphi^{-1}$ using the above discussion shows that Theorem 7.5 is equivalent to the following claim.
Claim.
Denote the rack structure of $\operatorname{Conj}X$ by $t$. Then for all functions $\rho^{\prime}:X\to X$, $\rho^{\prime}$ is a good involution of $\operatorname{Conj}X$ if and only if there exists a function $\overline{\alpha}:X\to Z(H)$ such that
$\rho^{\prime}(g,1)=\overline{\alpha}(g,1)(g,1),\qquad\overline{\alpha}t_{(g,1)}=\overline{\alpha}=\overline{\alpha}\rho^{\prime}$
for all $(g,1)\in X$.
By Proposition 7.4, all elements of $X$ and $Z(H)$ are involutions in $H$. Therefore, Theorem 6.7 provides the claim. ∎
@@
Let $\mathcal{O}:=G/\operatorname{Inn}(\operatorname{Core}G)$ be the set of orbits of $G$ under the action of $\operatorname{Inn}(\operatorname{Core}G)$. Then Theorem 7.5 states that every good involution $\rho$ of $\operatorname{Core}G$ factors through the natural projection $\pi:G\twoheadrightarrow\mathcal{O}$ and a unique function $\alpha^{*}:\mathcal{O}\to A$, and $\alpha^{*}$ completely determines $\rho$.
@@ 7.2.1. Corollaries
In light of Remark 7.6, we immediately obtain the following special cases of Theorem 7.5.
@@
$|A|\leq|{\operatorname{Good}(\operatorname{Core}G)}|\leq|A|^{|\mathcal{O}|}.$
@@
The lower bound in Corollary 7.7 is obtained from the constant functions $\alpha^{*}:\mathcal{O}\to A$; cf. Example 3.13. The upper bound is a consequence of Remark 7.6. ∎
@@
Suppose that $G$ is generated by the squares of its elements; that is, $G=\langle g^{2}\mid g\in G\rangle$. Then good involutions $\rho$ of $\operatorname{Core}G$ are precisely the multiplication maps $g\mapsto zg$ with $z\in A$. In particular, $\operatorname{Good}(\operatorname{Core}G)$ is in bijection with $A$.
@@
If $Z(G)$ is $2$-torsionless, then the only good involution of $\operatorname{Core}G$ is the identity map $\operatorname{id}_{G}$.
@@
Corollary 7.10 can also be obtained from Example 2.7 and Proposition 5.5.
@@
If $G$ is abelian, then Theorem 7.5 and Corollary 7.9 specialize to characterizations of good involutions of Takasaki kei $T(G)$. In particular, Corollary 7.9 applies to all Takasaki kei of $2$-divisible abelian groups; cf. [flat]. Moreover, Corollary 7.10 vastly generalizes a result of Kamada and Oshiro [symm-quandles-2]*Thm. 3.2 for dihedral quandles of odd order; cf. Example 5.6.
@@ 8. Algorithms
As applications of Theorems 6.7 and 7.5, we describe procedures for computing $\operatorname{Good}R$ when $R$ is a conjugation subquandle $\operatorname{Conj}X$ or a core quandle $\operatorname{Core}G$. We also discuss computational findings for groups up to order $23$.
@@ 8.1. How to compute $\operatorname{Good}(\operatorname{Conj}X)$
Let $n\geq 4$ be an even integer, and consider the dihedral group $D_{n}$ of order $2n$:
@@
$\rho(g):=(\zeta(g)g^{-1})^{-1}=\zeta(g)g^{-1}=\begin{cases}g&\text{if }g\in Z(D_{n}),\\ r^{n/2-k}&\text{if }g=r^{k}\text{ where }1\leq k<n\text{ and }k\neq\frac{n}{2},\\ sr^{n/2+k}&\text{if }g=sr^{k}\text{ for some }0\leq k<n.\end{cases}$
@@
Combined with Corollary 6.5 (or Proposition 5.3), Example 8.3 yields the following.
@@ 8.2. How to compute $\operatorname{Good}(\operatorname{Core}G)$
Let $G$ be a group of exponent greater than $2$, so that $\operatorname{Core}G$ is a nontrivial quandle. Similarly to the procedure for $\operatorname{Conj}X$, the following procedure constructs all good involutions $\rho$ of $\operatorname{Core}G$. The procedure is based on Theorem 7.5; cf. Remark 7.6.
@@
Let $\mathcal{O}$ be the set of orbits of $G$ under the action of $\operatorname{Inn}(\operatorname{Core}G)$, and let $\pi:G\twoheadrightarrow\mathcal{O}$ be the natural projection. Also, let $A$ be the $2$-torsion subgroup of the center $Z(G)$.
@@
    First, pick any function $\alpha^{*}:\mathcal{O}\to A$, and define functions $\alpha:G\to A$ and $\rho:G\to G$ by
@@
Given an integer $n\geq 2$, let $\operatorname{Dic}_{n}$ denote the dicyclic group of order $4n$:
@@
As in Algorithm 8.5, let $\pi:\operatorname{Dic}_{n}\twoheadrightarrow\mathcal{O}$ be the natural projection, and note that $A$ is the entire center $A=\{1,x^{2}\}$ of $\operatorname{Dic}_{n}$. On the other hand, routine calculations show that $\mathcal{O}$ contains exactly $(n+3)/2$ orbits:
  • one orbit $\pi(x)=\{x,x^{3}\}$ of size $2$,
  • one orbit $\pi(1)=\{a^{m}\mid 1\leq m\leq 2n\}$ of size $2n$, and
  • the following $(n-1)/2$ orbits of size $4$:
@@
In particular, the lower and upper bounds in Corollary 7.7 equal $2$ and $2^{(n+3)/2}$, respectively.
Evidently, each orbit in $\mathcal{O}$ is closed under multiplication by elements of $A$. That is,
$\pi(zg)=\pi(g)$
for all $z\in A$ and $g\in\operatorname{Dic}_{n}$. In particular, given any function $\alpha^{*}:\mathcal{O}\to A$, letting $g\in G$ be arbitrary and taking $z:=\alpha^{*}\pi(g)$ yields
@@
Since $\alpha^{*}:\mathcal{O}\to A$ was an arbitrary function, Theorem 7.5 states that the $2^{(n+3)/2}$ possible functions $\alpha^{*}$ collectively define $2^{(n+3)/2}$ distinct good involutions $\rho$ of $\operatorname{Core}D_{n}$, and no other good involutions exist. This completes the proof. ∎
@@
If $n\geq 4$ is a multiple of $4$, then the dihedral quandle $R_{n}=T(\mathbb{Z}/n\mathbb{Z})$ has exactly four good involutions. In particular, $\mathbb{Z}/n\mathbb{Z}$ exceeds the lower bound and attains the upper bound in Corollary 7.7.

### v5

@@ 7.1. Setup for the proof
To prove Theorem 1.4, we apply Observation 3.9 and Theorem 6.7 to a certain subquandle $\operatorname{Conj}X$ of a conjugation quandle $\operatorname{Conj}K$ such that $\operatorname{Conj}X\cong\operatorname{Core}G$. To study $\operatorname{Conj}X$, we first consider two subgroups of $Z(G)$.
7.1.1. The subgroups $A$ and $T$
Throughout this section, let $A$ be the subset of $Z(G)$ consisting of elements $a\in Z(G)$ satisfying the following condition: There exist a positive integer $n\in\mathbb{Z}^{+}$ and elements $g_{1},\ldots,g_{2n}\in G$ such that
$a=g_{1}g_{2}^{-1}g_{3}g_{4}^{-1}\cdots g_{2n-1}g_{2n}^{-1}=g_{1}^{-1}g_{2}g_{3}^{-1}g_{4}\cdots g_{2n-1}^{-1}g_{2n}.$
Observation 7.1.
$A$ is a subgroup of $Z(G)$ and, hence, a normal subgroup of $G$.
Throughout this section, let $T$ be the subgroup of $Z(G)$ whose elements have order 111 or 222:
$T:=\{z\in Z(G)\mid z^{2}=1\}.$
Proposition 7.2.
$T$ is a subgroup of $A$, and $T=A$ if $G$ is abelian or $Z(G)=T$.
Proof.
Straightforward. ∎
As an aside, we pose the following question.
Problem 7.3.
Does there exist a group $G$ such that $T\neq A$?
7.1.2. Core quandles as conjugation subquandles
Let $K$ be the semidirect product
@@ 7.1.2. Core quandles as conjugation subquandles
where $\mathbb{Z}/2\mathbb{Z}$ acts on $G\times G$ by swapping coordinates. In 2021, Bergman [core]*(6.5) showed that the function $\varphi:G\to K$ defined by
(7.1) $\varphi(g):=(g,g^{-1},1)$
is a quandle embedding of $\operatorname{Core}G$ into $\operatorname{Conj}K$. Henceforth, let $X:=\varphi(G)$ denote the image of $\varphi$, and let $H:=\langle X\rangle$ be the subgroup of $K$ generated by $X$. To prove Theorem 1.4, it suffices to apply Theorem 6.7 to $\operatorname{Conj}X$ by Observation 3.9.
Observation 7.4.
Every element of $X$ has order 222 in $H$.
Observation 7.5.
For all $g\in G$, $H$ contains the element $(g,g^{-1},0)$.
7.1.3. Calculation of the center
To apply Theorem 6.7 to $\operatorname{Conj}X$, we first compute $Z(H)$.
@@ 7.1.3. Calculation of the center
Let $G$ be a group containing an element $z$ such that $zgz^{-1}=g^{-1}$ for all $g\in G$. Then $G$ is an elementary abelian 222-group.
@@
Lemma 7.7.
Let $a\in Z(G)$. Then $(a,a,0)\in H$ if and only if $a\in A$.
Proof.
[…]
$(a,a,0)=\prod^{2n}_{i=1}(g_{i},g_{i}^{-1},1)=\prod^{2n-1}_{i\text{ odd}}(g_{i}g_{i+1}^{-1},g_{i}^{-1}g_{i+1},0)=(g_{1}g_{2}^{-1}\cdots g_{2n-1}g_{2n}^{-1},g_{1}^{-1}g_{2}\cdots g_{2n-1}^{-1}g_{2n},0)$
in $H$. Since $a\in Z(G)$, the equality of the first two coordinates shows that $a\in A$. ∎
@@
If $a\in A$, then $(a,a,0)\in H$ by Lemma 7.7, and verifying that $(a,a,0)\in Z(H)$ is straightforward.
Conversely, let $(a,b,i)\in Z(H)$ be arbitrary. If $i=1$, then Observation 7.5 shows that
$(g,g^{-1},0)=(a,b,1)(g,g^{-1},0)(a,b,1)^{-1}=(ag^{-1}a^{-1},bgb^{-1},0)$
for all $g\in G$. It follows from Lemma 7.6 that $G$ is an elementary abelian 222-group, which contradicts the assumption we made at the beginning of this section.
Thus, we must have $i=0$. By Lemma 7.7, it now suffices to show that $a\in Z(G)$ and $a=b$. Indeed,
$(g,g^{-1},0)=(a,b,0)(g,g^{-1},0)(a,b,0)^{-1}=(aga^{-1},bg^{-1}b^{-1},0)$
for all $g\in G$, so $a\in Z(G)$. On the other hand,
$(1,1,1)=(a,b,0)(1,1,1)(a,b,0)^{-1}=(ab^{-1},ba^{-1},1),$
so $a=b$. ∎
7.1.4. Preliminary versions of Theorem 1.4
We state two preliminary versions of Theorem 1.4. Denote the rack structure of $\operatorname{Conj}X$ by $t$.
Proposition 7.9.
For all functions $\rho^{\prime}:X\to X$, $\rho^{\prime}$ is a good involution of $\operatorname{Conj}X$ if and only if there exists a function $\zeta:X\to Z(H)$ such that
$\rho^{\prime}(g,g^{-1},1)=\zeta(g,g^{-1},1)(g,g^{-1},1),\qquad\zeta t_{(g,g^{-1},1)}=\zeta=\zeta\rho^{\prime}$
for all $(g,g^{-1},1)\in X$.
Proof.
This follows from combining Theorem 6.7 with Observation 7.4. ∎
In the following, denote the rack structure of $\operatorname{Core}G$ by $s$.
Proposition 7.10.
@@
We will show that the claim is equivalent to Proposition 7.9. Given any function $\alpha:G\to A$, define a function $\mathcal{F}(\alpha):X\to H$ by
$\mathcal{F}(\alpha)(g,g^{-1},1):=(\alpha(g),\alpha(g),0).$
By Proposition 7.8, the image of $\mathcal{F}(\alpha)$ lies in $Z(H)$. Conversely, given any function $\zeta:X\to Z(H)$, define a function $\mathcal{F}^{-1}(\zeta):G\to G$ by sending $g\in G$ to the first coordinate of $\zeta(g,g^{-1},1)\in Z(H)$. By Proposition 7.8, the image of $\mathcal{F}^{-1}(\zeta)$ lies in $A$. Hence, it is straightforward to verify that the assignments $\mathcal{F}$ and $\mathcal{F}^{-1}$ are mutually inverse.
Now, recall that the map $\varphi$ from (7.1) is a quandle isomorphism from $\operatorname{Core}G$ to $\operatorname{Conj}X$. The equivalence of the claim with Proposition 7.9 therefore follows from Observation 3.9 and the following straightforward calculations: For all elements $g\in G$ and functions $\rho:G\to G$ and $\rho^{\prime}:X\to X$ respectively satisfying the conditions of the claim and Proposition 7.9, we have
$\varphi\rho\varphi^{-1}(g,g^{-1},1)=\mathcal{F}(\alpha)(g,g^{-1},1)(g,g^{-1},1),\qquad\varphi^{-1}\rho^{\prime}\varphi(g)=\mathcal{F}^{-1}(\zeta)(g)g,$
so that
$\mathcal{F}(\alpha s_{g})=\mathcal{F}(\alpha)t_{(g,g^{-1},1)},\qquad\mathcal{F}(\alpha\rho)=\mathcal{F}(\alpha)\varphi\rho\varphi^{-1}$
and, similarly,
$\mathcal{F}^{-1}(\zeta t_{(g,g^{-1},1)})=\mathcal{F}^{-1}(\zeta)s_{g},\qquad\mathcal{F}^{-1}(\zeta\rho^{\prime})=\mathcal{F}^{-1}(\zeta)\varphi^{-1}\rho^{\prime}\varphi.$
7.2. Main results
We now prove Theorem 1.4. Preserve the notation from the previous subsection, so that $T$ denotes the subgroup $T=\{z\in Z(G)\mid z^{2}=1\}$.
Theorem 7.11.
Proposition 7.10 holds when $A$ is replaced with $T$.
Proof.
If $\alpha:G\to T$ is a function satisfying the conditions of Proposition 7.10, then $\alpha$ induces a good involution $\rho$ by Propositions 7.2 and 7.10.
Conversely, let $\rho$ be a good involution of $\operatorname{Core}G$. Then $\rho$ is induced by a unique function ${\alpha:G\to A}$ satisfying the conditions of Proposition 7.10, and for all $g\in G$, we have
$s_{\rho(g)}=s_{g}^{-1}=s_{g}$
because $\operatorname{Core}G$ is involutory. Hence, [core2]*Lem. 4.1 states that
$T\ni g^{-1}\rho(g)=g^{-1}\alpha(g)g=\alpha(g),$
where we have used the fact that $\alpha(g)\in Z(G)$. Since $g\in G$ was arbitrary, the image of $\alpha$ lies in $T$, as desired. ∎
Example 7.12.
The constant functions $\alpha:G\to T$ recover Example 3.14.
@@ 7.2. Main results
Let $\mathcal{O}:=G/\operatorname{Inn}(\operatorname{Core}G)$ be the set of orbits of $G$ under the action of $\operatorname{Inn}(\operatorname{Core}G)$. Then Theorem 7.11 states that every good involution $\rho$ of $\operatorname{Core}G$ factors through the natural projection $\pi:G\twoheadrightarrow\mathcal{O}$ and a unique function $\alpha^{*}:\mathcal{O}\to T$, and $\alpha^{*}$ completely determines $\rho$.
@@ 7.2.1. Corollaries
Using Remark 7.13, we obtain the following special cases of Theorem 7.11.
@@
$|T|\leq|{\operatorname{Good}(\operatorname{Core}G)}|\leq|T|^{|\mathcal{O}|}.$
@@
The lower bound in Corollary 7.14 is obtained from the constant functions $\alpha^{*}:\mathcal{O}\to T$; cf. Example 7.12. The upper bound is a consequence of Remark 7.13. ∎
@@
Suppose that $G$ is generated by the squares of its elements; that is, $G=\langle g^{2}\mid g\in G\rangle$. Then good involutions $\rho$ of $\operatorname{Core}G$ are precisely the multiplication maps $g\mapsto tg$ with $t\in T$. In particular, $\operatorname{Good}(\operatorname{Core}G)$ is in bijection with $T$.
@@
If $Z(G)$ is 222-torsionless, then the only good involution of $\operatorname{Core}G$ is the identity map $\operatorname{id}_{G}$.
@@
Corollary 7.17 can also be obtained from Example 2.7 and Corollary 5.5.
@@
If $G$ is abelian, then Theorem 7.11 and Corollary 7.16 specialize to characterizations of good involutions of Takasaki kei $T(G)$. In particular, Corollary 7.16 applies to all Takasaki kei of 222-divisible abelian groups; cf. [flat]. Moreover, Corollary 7.17 vastly generalizes a result of Kamada and Oshiro [symm-quandles-2]*Thm. 3.2 for dihedral quandles of odd order; cf. Example 5.6.
@@ 8. Algorithms
As applications of Theorems 6.7 and 7.11, we describe procedures for computing $\operatorname{Good}R$ when $R$ is a conjugation subquandle $\operatorname{Conj}X$ or a core quandle $\operatorname{Core}G$. We also discuss computational findings for groups up to order 232323.
@@ 8.1. How to compute $\operatorname{Good}(\operatorname{Conj}X)$
[…]
@@
$\rho(g):=(\zeta(g)g^{-1})^{-1}=\begin{cases}g&\text{if }g\in Z(D_{n}),\\ r^{n/2-k}&\text{if }g=r^{k}\text{ where }1\leq k<n\text{ and }k\neq\frac{n}{2},\\ sr^{n/2+k}&\text{if }g=sr^{k}\text{ for some }0\leq k<n.\end{cases}$
@@
Combined with Corollary 6.5 (or Corollary 5.3), Example 8.3 yields the following.
@@ 8.2. How to compute $\operatorname{Good}(\operatorname{Core}G)$
Let $G$ be a group of exponent greater than 222, so that $\operatorname{Core}G$ is a nontrivial quandle. Similarly to the procedure for $\operatorname{Conj}X$, the following procedure constructs all good involutions $\rho$ of $\operatorname{Core}G$. The procedure is based on Theorem 7.11; cf. Remark 7.13.
@@
Let $\mathcal{O}$ be the set of orbits of $G$ under the action of $\operatorname{Inn}(\operatorname{Core}G)$, and let $\pi:G\twoheadrightarrow\mathcal{O}$ be the natural projection. Also, let $T:=\{z\in Z(G)\mid z^{2}=1\}$.
@@
    First, pick any function $\alpha^{*}:\mathcal{O}\to T$, and define functions $\alpha:G\to T$ and $\rho:G\to G$ by
@@
[…]
@@
[…]
  • one orbit $\pi(x)=\{x,x^{3}\}$ of size 222,
[…]
  • the following $(n-1)/2$ orbits of size 444:
@@
In particular, the lower and upper bounds in Corollary 7.14 equal 222 and $2^{(n+3)/2}$, respectively.
Evidently, each orbit in $\mathcal{O}$ is closed under multiplication by elements of $T$. That is,
$\pi(tg)=\pi(g)$
for all $t\in T$ and $g\in\operatorname{Dic}_{n}$. In particular, given any function $\alpha^{*}:\mathcal{O}\to T$, letting $g\in G$ be arbitrary and taking $t:=\alpha^{*}\pi(g)$ yields
@@
Since $\alpha^{*}:\mathcal{O}\to T$ was an arbitrary function, Theorem 7.11 states that the $2^{(n+3)/2}$ possible functions $\alpha^{*}$ collectively define $2^{(n+3)/2}$ distinct good involutions $\rho$ of $\operatorname{Core}D_{n}$, and no other good involutions exist. This completes the proof. ∎
@@
If $n\geq 4$ is a multiple of 444, then the dihedral quandle $R_{n}=T(\mathbb{Z}/n\mathbb{Z})$ has exactly four good involutions. In particular, $\mathbb{Z}/n\mathbb{Z}$ exceeds the lower bound and attains the upper bound in Corollary 7.14.